What is the total amount you will have to repay for your $41,000 student loan if the interest rate is 4% APR and you pay equal monthly payments over the 10 year life of the loan?
Answer #1

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

r = 4%/12 = 0.33% (monthly), n = 10 * 12 = 120 months, PV = $41,000

41000 =P*1-(1+0.0033) – 120 0.0033

41000 = P * 98.7702

P = $415.11